OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! A tool box has a volume of x^3 + 8x^2 + 11x – 20 cm^3 and the height is x + 5 cm. Find the polynomial that would represent the area of the bottom of the tool box? Explain your reasoning. @jim_thompson5910 @satellite73 @thomaster @Mertsj

OpenStudy (magbak):

@Mertsj @jim_thompson5910 @satellite73 @thomaster

OpenStudy (anonymous):

divide

OpenStudy (magbak):

How.

OpenStudy (anonymous):

\[( x^3 + 8x^2 + 11x – 20 )\div (x+5)\]

OpenStudy (anonymous):

1) long division (ick) 2) synthetic division (very snappy) 3) thinking (it aint illegal yet)

OpenStudy (magbak):

Preferably synthetic division. Can you explain it please.

OpenStudy (anonymous):

i can explain it but it is an amazing pita to write it here

OpenStudy (anonymous):

how about we think?

OpenStudy (magbak):

Ok

OpenStudy (anonymous):

you are dividing a polynomial of degree 3 by a polynomial of degree 1, leaving a polynomial of degree 2, i.e. a quadratic, and by the set up of the question is it pretty obvious it will go in there evenly so we start

OpenStudy (anonymous):

\[ x^3 + 8x^2 + 11x – 20 =(x+5)(ax^2+bx+c)\] now it should be rather obvious that i was being silly writing \(a\) when we know \(a=1\) since that is the only way we are going to get \(x^3\) \[ x^3 + 8x^2 + 11x – 20 =(x+5)(x^2+bx+c)\]

OpenStudy (anonymous):

also we know \(c\) right away , since the constant is \(-20\) and we have a \(+5\) in the first term, making \(c=-4\)

OpenStudy (anonymous):

\[x^3 + 8x^2 + 11x – 20 =(x+5)(x^2+bx-4)\] so both \(a\) and \(c\) are clear from the start, i was only writing that other stuff to show what i was doing the only think part is finding \(b\)

OpenStudy (magbak):

OK I am sorry can you slow do my thinking process is a little slow please.

OpenStudy (anonymous):

ok really you need only look at this line \[x^3 + 8x^2 + 11x – 20 =(x+5)(x^2+bx-4)\] and see that it is clear that the first term on the quadratic must be \(x^2\) and the constant must be \(-4\) when that is clear let me know and we can find \(b\)

OpenStudy (magbak):

where do you get b from.

OpenStudy (anonymous):

we didn't get it yet, i was slowing down to make sure it is clear to you that the quadratic must be \(x^2+bx-4\) so all we need is \(b\) to complete it is that clear? i.e. is it clear that it must start with \(x^2\) and end with \(-4\) ?

OpenStudy (magbak):

But the original equation does not even have a B. That was what I was asking.

OpenStudy (anonymous):

oh \(b\) is just some number we need to find that is all the work , which is not hard at all

OpenStudy (anonymous):

i just used \(b\) because frequently we write a quadratic as \(ax^2+bx+c\) so i used \(b\) for the coefficient of the \(x\) term

OpenStudy (magbak):

I have a question is a volume consisting of lxwxh so we only need the lxw no height right.

OpenStudy (anonymous):

forget the verbiage of the question, which is nonsense , just trying to make it sound like a word problem your job is to divide, that is all

OpenStudy (anonymous):

the question is really asking you to divide \[( x^3 + 8x^2 + 11x – 20 )\div (x+5)\]

OpenStudy (magbak):

Ok so all we do is divide So their is no 3 monomials right? That is the last stupid thing I will ask.

OpenStudy (magbak):

Oh ok.

OpenStudy (anonymous):

no , it asks for the area which will be your quadratic

OpenStudy (magbak):

OK givem e a sek to divide.

OpenStudy (anonymous):

there might be three monomials if your quadratic factors, but it doesn't ask for them

OpenStudy (anonymous):

lol i though that is what we were doing

OpenStudy (magbak):

So it is x^2 +3x -4 with a remainder of 0.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where did you get the 3 from? it is right, i am just asking

OpenStudy (magbak):

ok so just if I am not asking to much can you please word the problem answer for me.

OpenStudy (magbak):

When I divided that is what came up.

OpenStudy (anonymous):

i guess i meant by what method, but nvm

OpenStudy (anonymous):

Find the polynomial that would represent the area of the bottom of the tool box? Explain your reasoning.

OpenStudy (anonymous):

the volume is length times width times height since we know the volume is \(x^3+8x^2-11x-20\) and the height is \(x+5\) we know that length times width, the area of the base, is \((x^3+8x^2+11x-20)\div (x+5)\)

OpenStudy (magbak):

Is that the complete answer? @satellite73

OpenStudy (magbak):

Thank you very much @satellite73 I have only 3 more questions I hope you can help me with. Thank you again.